A rectangle has a perimeter of 34 inches. The length is 7 inches longer than the width.What is the area?

Answer by Boreal(15235)

width=x
length=x+7
product is 34
x(x+7)=34
x^2+7x-34=0
x=(1/2)(-7+/-sqrt(49+136)); sqrt 185=13.60
x=(1/2)(6.60), taking positive root only
x=3.30 in width
x+7=10.30 in length
Their product is 33.99 sq in